OpenStudy (anonymous):

A soccer ball is kicked toward the goal. The height of the ball is modeled by the function h(t) = -16t2 + 48t where t equals the time in seconds and h(t) represents the height of the ball at time t seconds. What is the axis of symmetry, and what does it represent? t = 1.5; it takes 1.5 seconds to reach the maximum height and 3 seconds to fall back to the ground t = 1.5; it takes 1.5 seconds to reach the maximum height and 1.5 seconds to fall back to the ground t = 3; it takes 3 seconds to reach the maximum height and 3 seconds to fall back to the ground t = 3; it takes 3 second

OpenStudy (anonymous):

to reach the maximum height and 4 seconds to fall back down to the ground

OpenStudy (imstuck):

In order to find the axis of symmetry, you need to set the equation up for graphing in the form\[(x-h)^{2}=y-k\]

OpenStudy (imstuck):

To do that you need to complete the square on the x terms. Hold on and I will post.

OpenStudy (anonymous):

what do u mean? well the 16t^2 is actually squared already

OpenStudy (imstuck):

I know but it doesn't matter. Getting it into this form puts it just how you need it, like this:\[y-36=-16(t-\frac{ 3 }{ 2 })^{2}\]This tells you that the axis of symmetry is the line x = 3/2.

OpenStudy (anonymous):

Ok, but i don't understand how you get the 3/2?

OpenStudy (imstuck):

The axis of symmetry represents the point where the ball is at its highest point.

OpenStudy (imstuck):

I'll show you how to complete the square to get that.

OpenStudy (anonymous):

ok thank you.

OpenStudy (imstuck):